# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

### v4

@@ 5 Experiments
We conduct extensive experiments to better understand the LT3D  problem and gain insights by validating our techniques described in Sec. 4. Specifically, we aim to answer the following questions: [^1] Answers: yes, yes, yes, yes.
@@
    Does multi-modal late-fusion help detect rare classes?
We compare our multi-modal late-fusion (MMLF) approaches (Fig. 5 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig. 3. Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table I).
Datasets. To explore LT3D, we use the well established nuScenes [6] and Argoverse 2.0 (AV2) [4] datasets. Other AV datasets such as KITTI [2] and Waymo [5] do not support the study of LT3D as they only annotate three common classes. nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions. To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [44]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [74] and ImageNet-LT (IF=1,000) [75]. nuScenes arranges classes in a semantic hierarchy (Fig. 4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (please refer to the supplement for details). Following prior work, we use official training and validation sets. We focus on nuScenes in the main paper and AV2 in the supplement. Our primary conclusions hold for both datasets.
Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [11]) and 3D LiDAR detections (from CenterPoint [12]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
TABLE I: Benchmarking results on nuScenes. We denote the RGB and LiDAR modalities by C and L, respectively. Our MMLF fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic ensembling. MMLF performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [35], DeepInteraction [59], and CMT [10]. Importantly, our MMLF approach nearly doubles the detection performance achieved by prior work on classes with few examples!
@@
FCOS3D [13] C 20.9 39.0 23.3 2.9
BEVFormer [23] C 27.3 52.3 31.6 1.4
PolarFormer [25] C 28.0 54.0 31.6 2.2
CenterPoint [12] L 40.4 77.1 45.1 4.3
TransFusion-L [34] L 38.5 68.5 42.8 8.4
BEVFusion-L [35] L 42.5 72.5 48.0 10.6
CMT-L [10] L 34.7 73.4 35.9 1.1
CLOCS [38] C+L 40.0 68.2 45.7 10.0
TransFusion [34] C+L 39.8 73.9 41.2 9.8
BEVFusion [35] C+L 45.5 75.5 52.0 12.8
DeepInteraction [59] C+L 43.7 76.2 51.1 7.9
CMT [10] C+L 44.4 79.9 53.0 4.8
CenterPoint [12] + RCNN [76] C+L 34.0 64.8 37.5 4.3
MMF(CenterPoint, FCOS3D) (Ours) C+L 43.6 77.1 49.0 9.4
MMLF(CenterPoint, DINO) (Ours) C+L 51.4 77.9 59.4 20.0
5.1 State-of-the-Art Comparison on nuScenes
Table I compares our MMF and MMLF approaches with prior work on nuScenes. Fig. 9 presents qualitative results. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
CenterPoint [12], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 4.3 mAP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on classes with many examples. We posit that limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance. In contrast, BEVFusion [23], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [76] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4 mAP, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy. This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier. Our multimodal filtering method MMF(CenterPoint, FCOS3D) keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions. This simple baseline achieves 9.4 mAP on classes with few examples. Lastly, we re-implement CLOCS [38] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections. However, CLOCs performs worse than our MMF baseline. By carefully considering design choices outlined in Fig. 3, MMLF(CenterPoint, DINO) improves over MMF(CenterPoint, FCOS3D) by 7.8 mAP!
5.2 Ablation Studies and Empirical Analysis
We conduct extensive ablations to validate our proposed MMLF approach. We provide additional results in the supplement.
TABLE II: Ablation on Hierarchy Loss and Multi-Modal Filtering (MMF). Training with the semantic hierarchy improves both PointPillars [17] and CenterPoint [12] for LT3D by $>$1% mAP over All classes. Moreover, MMF yields 4$\sim$11 mAP improvement on Medium and Few classes, showing the benefit of late fusion!
Method Modality All Many Medium Few
FCOS3D [13] C 20.9 39.0 23.3 2.9
PointPillars [17] L 30.0 64.2 28.4 3.4
@@ 5.2 Ablation Studies and Empirical Analysis
+ MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
5.2.1 Analysis on Training with Semantic Hierarchy
We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors (PointPillars [17] and CenterPoint [12]) to jointly predict class labels at different levels of the semantic hierarchy. For example, we modify the detector to additionally classify stroller as pedestrian and object. The semantic hierarchy naturally groups classes based on shared attributes and may have complementary features, helping train a detector that allows feature sharing across common and rare classes and leads to improved performance. Table II highlights that this approach (“+ Hierarchy Loss”) effectively improves LT3D accuracy by $>$1 mAP over all classes.
TABLE III: Fusing uni-modal detections in 3D vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchy loss) with 2D RGB and 3D RGB detections in 3D versus on the image plane. We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF). Following [70], we inflate 2D detections to 3D using LiDAR points within the box frustum. We project 3D detections to the image plane using provided sensor extrinsics. Results show that matching 3D RGB detections in 3D and on the image plane yields similar results. Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D. In contrast, filtering LiDAR-based detections using 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP. This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors. We further verify this in the supplement.
@@ 5.2.1 Analysis on Training with Semantic Hierarchy
+ FCOS3D [13] 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
+ BEVFormer [23] 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
+ PolarFormer [25] 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
+ YOLOV7 [58] 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
+ DINO [11] 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
TABLE IV: Progressively adding training components improves LT3D performance, confirming that 2D RGB detectors are better suited for multi-modal late-fusion, matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and score calibration prior and probabilistic fusion improves performance. Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO. In addition, training with external data further improves rare class accuracy by 4.2 mAP.
@@
 + Probabilistic Ensembling +0.9 51.4 77.9 59.4 20.0
Figure 10: Both our MMLF method (columns 1-3) and TransFusion [34] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult.
5.2.2 Analysis on Multi-Modal Late-Fusion
We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table III. We further examine the impact of using additional data and study different fusion strategies in Table IV.
How do we incorporate RGB information? Table II demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D. We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table III. Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
How do we match detections from uni-modal detectors? MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections. This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table III. Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, notably achieving marginally lower accuracy than the LiDAR-only baseline.
Table III shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10 mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that MMLF performs better with 2D RGB detectors than 3D RGB detectors.
How do we fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7 mAP averaged over all classes (Table IV). Once calibrated, we probabilistically fuse matched detections to generate the final set of detections. Table IV shows that probabilistic fusion improves performance by 0.9 mAP.
TABLE V: Diagnosis using the mAPH metric on selected classes. We analyze CenterPoint [12] and CMT [10], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [11]). Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
@@ 5.2.2 Analysis on Multi-Modal Late-Fusion
CenterPoint LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
(Group-Free) LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
@@
CenterPoint LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
(Group-Free) LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
+ Hierarchy Loss LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
(Group-Free) LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
+ Hier. + MMLF LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
@@
CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
(LiDAR-only) LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
+ MMLF LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3

### v5

@@ 6 Experiments
We conduct extensive experiments to better understand the LT3D problem and gain insights by validating our techniques described in Sec. 4. We aim to answer the following questions: [^1] Answers: yes, yes, yes, yes, yes.
@@
    Does our multimodal late-fusion strategy resoundingly outperform previous methods?
  5. 5.
    Does our multi-modal late-fusion help detect rare classes?
We compare our Multi-Modal Filter (MMF) and Multi-Modal Late-Fusion (MMLF) approaches (Fig. 6 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig. 3. Regarding notation, MMF(X, Y) and MMLF(X, Y) mean that they fuse detections from the LiDAR-based detector X and the RGB-based detector Y. Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table 1). We provide implementation details in Appendix A.
Datasets. To explore LT3D, we use the well-established nuScenes [5] and Argoverse 2.0 (AV2) [73] datasets. Other AV datasets such as KITTI [21] and Waymo [64] do not support the study of LT3D as they annotate only three common classes. In contrast, nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions. To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [10]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [68] and ImageNet-LT (IF=1,000) [38]. nuScenes arranges classes in a semantic hierarchy (Fig. 4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (refer to Appendix B for details). Following prior work, we use the official training set for training and the validation set for benchmarking. As our primary conclusions hold for both datasets, we report results on nuScenes in the main paper and provide results on AV2 in Appendix B.
Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by fusing 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child. Refer to Appendix O for 3D visualizations.
@@
FCOS3D C 20.9 39.0 23.3 2.9
PointPillars (PP) L 30.0 64.2 28.4 3.4
@@
 + MMF(CP, FCOS3D) C+L 43.6 77.1 49.0 9.4
[…]
6.1 Benchmarking Results
Table 1 compares our MMF and MMLF with prior work on nuScenes. Fig. 9 displays qualitative results. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes. We provide breakdown analysis per class in Appendix H and show results on the standard 10-class benchmark in Appendix K.
CenterPoint [82], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving only 4.3 mAP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Yet, it is unexpected that the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on many classes. We posit that the limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance (Appendix D). In contrast, BEVFusion [43], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
@@ 6.1 Benchmarking Results
+ FCOS3D 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
+ BEVFormer 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
+ PolarFormer 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
+ YOLOV7 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
+ DINO 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
Table 3: Fusing uni-modal detections in 3D vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with our hierarchy loss) with 2D RGB vs. and 3D RGB detections, in 3D vs. on the image plane. We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF). When fusing in 3D, we follow [72] that inflates 2D detections to 3D using LiDAR points within their box frustums. When fusing on the 2D image plane, we project 3D detections to the image plane using provided sensor extrinsics. When fusing 3D RGB detections, similar results are achieved regardless matching in 3D or on the 2D image plane. Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D. Yet, late-fusion between LiDAR-based 3D detections and RGB-based 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by >>10 mAP. Refer to Appendix I for more results.
@@
 + Probabilistic Fusion +0.9 51.4 77.9 59.4 20.0
Table 4: Adding our developed components progressively improves LT3D performance, confirming that (1) 2D RGB detectors are better suited for our MMLF, (2) matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and (3) score calibration prior and probabilistic fusion improves performance. A majority of performance improvements is attributed to fusing detections of the 2D RGB detector DINO on the 2D image plane. Moreover, training with external data notably improves rare class accuracy by 4.2 mAP.
Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [18] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4 mAP, likely because it learns classifiers on cropped regions that does not exploit contextual information useful for recognition. This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier. Our multimodal filtering method MMF keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions. This simple baseline achieves 9.4 mAP on classes with few examples. Lastly, we re-implement CLOCS [49] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections. However, CLOCs performs worse than our MMF baseline. By carefully considering design choices outlined in Fig. 3, MMLF improves over MMF by 7.8 mAP!
6.2 Ablation Study
We conduct extensive ablations to validate our proposed MMLF approach. In appendices, we provide additional results on the impact of calibration noise (Appendix N), detector recall (Appendix M) and performance breakdowns by object visibility and distance from ego-vehicle (Appendix L).
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller $P$
@@ 6.2 Ablation Study
CenterPoint (w/ our group-free head) LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
@@
CenterPoint (w/ our group-free head) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
+ Hierarchy Loss LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint (w/ our group-free head) LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
+ Hier. + MMLF LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
@@
CMT-L (LiDAR-only) LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
+ MMLF LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
Table 5: Diagnosis using the mAPH metric on selected classes. We analyze the LiDAR-based 3D detector CenterPoint [82] and mutli-modal detector CMT [80], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [88]). Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable ($P$). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
6.2.1 Analysis on Training with Semantic Hierarchy
We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors to jointly predict class labels at different levels of the hierarchy. For example, we modify detectors PointPillars [41] and CenterPoint [82] to additionally classify stroller as pedestrian and object. As the hierarchy naturally groups classes based on shared attributes, training with it helps learn complementary features across levels that are shared across common and rare classes. Table 2 highlights that this approach (“+ Hierarchy Loss”) achieves improvements on all class by >>1 mAP.
6.2.2 Analysis on Multi-Modal Late-Fusion
We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table 3. We further examine the impact of using additional data and study different fusion strategies in Table 4.
How to incorporate RGB information? Table 2 (+ MMF) demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D. We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table 3. Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
Figure 10: Both our MMLF method (columns 1-3) and TransFusion [2] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult. Refer to Appendix O for more visualizations.
How to match detections from uni-modal detectors? MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections. This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table 3. Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, achieving marginally lower accuracy than the LiDAR-only baseline.
Notably, projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by >>10 mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, due to the erroneous 3D predictions that affects 2D detection quality.
How to fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7 mAP averaged over all classes (Table 4). With calibration, our probabilistic fusion fuses matched detections to generate the final detections, achieving 0.9 mAP improvements (Table 4).